Hello, I have three questions for you.
1. The number of entries in a race decreased 20% from 530 entries in last years race. Find the new number of entries.
2. The number of accounts with a certain firm increases 30% from a level of 200 accounts. Find the new number of accounts.
3. The number of gas station in a city increasesfrom 120 to 150. Find the percent increase.

Found 2 solutions by Abbey, JimMarshall:
Answer by Abbey(339)

1.
If the number of entries decreased by 20%, then it is only 80% of last year:
530 * .80 = 424 entries
2. 30% of 200 is 60 (.3 * 200 = 60). The number of accounts now totals 260 accounts.
3. The increase in gas stations is 30. ( 150 - 120 = 30 ). Set this problem up as a proportionality
30/120 = x/100
and cross multiply
120x = 3000
Divide both sides by 120
x = 25 (%) increase

Answer by JimMarshall(82)

If the number of entries decreased by 20%, then it is only 80% of last year:
530 * .80 = 424 entries
2. 30% of 200 is 60 (.3 * 200 = 60). The number of accounts now totals 260 accounts.
3. The increase in gas stations is 30. ( 150 - 120 = 30 ). Set this problem up as a proportionality
30/120 = x/100
and cross multiply
120x = 3000
Divide both sides by 120
x = 25 (%) increase